On the opening night of the Broadway musical Miss Saigon, the Rhapsody Theater sold 800 total tickets. An advanced ticket costs $14.50 and a ticket at the door costs $22.00. In all. $16,640 was taken in. How many tickets were sold at he door?
Answer by Abbey(339)

let x = the number of advanced tickets sold
14.50x = the amount of money the advanced tickets made (the number of tickets * the price)
Let y = the number of door tickets sold
let 22y=the amount of money the door tickets made.
The total revenue is:
14.5x+22y=16640
The total # of tickets sold is:
x+y = 800
subtract x from both sides:
y=800-x
then substitute 800-x into the other equation:
14.5x+22y=16640
14.5x+22(800-x)=16640
clear the parentheses:
14.5x+17600-22x=16640
combine like terms; subtract 17600 from both sides:
-7.5x=16640-17600
-7.5x=-960
divide both sides by -7.5:
x=128
If 128 advanced tickets were sold, then 800-128=672 tickets were sold at the door
Check your answer:
128*14.5=1856
672*14784
1856+14784=16640